# Bidirectional Reward-Guided Diffusion for Real-World Image Super-Resolution

### v1

@@ 5.1 Experimental Settings
Datasets. We adopt a combination of images from DIV2K (Agustsson and Timofte, 2017), DIV8K (Gu et al., 2019), Flickr2K (Timofte et al., 2017), and the first 10K face images from FFHQ (Karras et al., 2019) during training. Flickr8K (Cui et al., 2024b) is used for real world low-resolution dataset. The degradation pipeline of RealESRGAN (Wang et al., 2021) is utilized to synthesize LR-HR training pairs with the same parameter configuration as baseline methods. Following DiT4SR (Duan et al., 2025), the resolutions are set to 128 × 128 and 512 × 512 for LR and HR images, respectively. Following ResShift (Yue et al., 2023), the resolutions are set to 64 × 64 and 256 × 256 for LR and HR images.
Since our method specifically focuses on Real-ISR task, we evaluate our model on four widely used real-world datasets, including DrealSR (Wei et al., 2020), RealSR (Cai et al., 2019), RealLR200 (Wu et al., 2024b), and RealLQ250 (Ai et al., 2024). All experiments are conducted with the scaling factor of ×4. DrealSR and RealSR respectively consist of 93 and 100 images. Following DiT4SR, center-cropping is adopted for these two datasets, and the resolution of LR images is set to 128 × 128.
Metrics. Following (Duan et al., 2025; Ai et al., 2024; Yu et al., 2024), PSNR and SSIM (Wang et al., 2004) are inadequate for measuring perceptual differences. Thus, most studies use the perceptual measurement LPIPS (Zhang et al., 2018a) for image fidelity. We use MUSIQ (Ke et al., 2021), MANIQA (Yang et al., 2022), ClipIQA (Wang et al., 2023), and LIQE (Zhang et al., 2023b) as non-reference metrics to measure image quality. In addition, we conduct a user study to comprehensively assess both fidelity and perceptual quality.
@@
Table 2: User study results on real-world datasets. The percentages denote the frequency with which DIT4SR w/Ours was preferred over each compared approach, for both realism and fidelity.
@@
5.2 Comparison with Other Methods
We compare our method with state-of-the-art RealISR methods, including diffusion-based approaches (i.e., ResShift (Yue et al., 2023), StableSR (Wang et al., 2024a), SeeSR (Wu et al., 2024b), DiffBIR (Lin et al., 2024), and SUPIR (Yu et al., 2024)), as well as diffusion-based methods with DiT architecture (i.e. DreamClear (Ai et al., 2024) and DiT4SR (Duan et al., 2025)). As shown in Table 1, our method achieves strong quantitative performance across all four real-world super-resolution benchmarks. Additional visual results, including the outputs of ResShift, are presented in the appendix. As illustrated in Figure 3, our approach produces more realistic and natural reconstructions in real-world scenarios, with clearer and more faithful details while effectively suppressing artifacts. These results indicate that our method effectively mitigates the impact of input distribution shift and improves perceptual quality on real-world low-resolution images. By leveraging bidirectional reward feedback learning and jointly training on synthetic and real low-resolution data, our method demonstrates superior performance and robustness.
5.3 Ablation Experiments
To further demonstrate the effectiveness of each component, we conduct the ablation study on RealSR with MUSIQ and LPIPS as evaluation metrics. All variants are trained using the same settings as the full model for fair comparison. Additional details are provided in the Appendix.
@@ 5.3.1 Effectiveness of Bidirectional Reward-Guided Diffusion
Setting MUSIQ $\uparrow$ LPIPS $\uparrow$
@@
4. Mixed forward and reverse: The full bidirectional optimization framework that jointly leverages synthetic paired data and real-world LR data.
5.3.2 dynamic weights
We investigate the impact of different distortion–perception weighting strategies across diffusion timesteps on model performance. Specifically, we compare several weighting schedules for combining distortion loss $\mathcal{L}_{\text{struct}}$ and reward-based loss $\mathcal{L}_{\text{reward}}$ in our bidirectional reward-guided SR framework: $\lambda(t)=\left(\frac{t}{T}\right)^{\gamma}$, which gradually shifting from distortion to perceptual reward, where $\gamma>0$ controls the decay rate.
Analysis. Table 4 summarizes the quantitative results, and Figure 5 visualizes the different weighting schedules. Our ablation study on power-law weighting with varying exponents reveals a critical trade-off (Blau and Michaeli, 2018; Sun et al., 2024c; Jinjin et al., 2020): stronger structural constraints at early timesteps reduces distortion but limits perceptual reward gains. Conversely, insufficient early-stage constraints lead to structural degradation, which triggers reward hacking. By identifying an optimal weighting scheme, we enhance perceptual preference while preserving structural integrity.
@@
Table 4: Ablation study on different distortion–perception weighting. We report perceptual quality measured by MUSIQ and LPIPS.
@@ 6 Conclusion
Appendix A Detailed Formulations of Diffusion-based Super-Resolution
@@
A.1 Diffusion Models Trained from Scratch for Super-Resolution
The first category directly trains a diffusion model for super-resolution using paired low-resolution (LR) and high-resolution (HR) data. Representative methods such as ResShift formulate the forward diffusion process as an interpolation between the HR image $x_{0}$, the LR-conditioned signal $x_{1}$, and Gaussian noise.
@@ A.1 Diffusion Models Trained from Scratch for Super-Resolution
A.2 Pretrained Diffusion Models with Conditional Adaptation
@@
A common approach is to introduce an external conditioning module, such as ControlNet, which injects LR structural information into intermediate layers of the frozen pretrained diffusion backbone. Alternatively, lightweight parameter-efficient tuning strategies, such as Low-Rank Adaptation (Hu et al., 2022) (LoRA), are employed to fine-tune a subset of model parameters while preserving the pretrained generative prior. Specifically, the forward diffusion process remains unchanged and is given by:
@@
where $x_{0}$ denotes the target high-resolution image. The low-resolution observation $y$ is not involved in the forward noising process but is provided as a conditioning signal to the denoising network. Formally, the reverse diffusion process can be written as:
@@
Appendix B Implementation details
B.1 Local Binary Pattern (LBP) Texture Features
@@
B.2 KDE of Cosine Similarity Distributions in Deep Feature Space
We further analyze the relationship between low-resolution and high-resolution (HR) image pairs in deep feature space by computing the cosine similarity between their feature embeddings extracted from a pretrained network. Kernel density estimation (KDE) is employed to estimate the distribution of cosine similarities across the dataset. Across different frequency bands, synthetic data exhibits a notable distribution expansion coupled with a significant domain shift, as shown in Figure 8. In the low-frequency band, synthetic LR–GT pairs maintain high cosine similarity concentrated near 1.0, closely matching real-world distributions and preserving global structures. However, as the frequency increases, the synthetic distribution becomes increasingly broader and shifts toward lower similarity values; specifically, the mid-frequency band reveals discrepancies in meso-scale textures, while the high-frequency band displays the most pronounced flattening and leftward shift. This trend suggests that while synthetic degradations cover a wider range of variations, they fail to accurately replicate the complex, fine-grained artifacts of real-world processes, identifying high-frequency components as the primary driver of the overall input distribution shift. Therefore, when performing reward feedback learning on real-world data, we primarily apply optimization at the final timestep. This is because, during early timesteps, the diffusion model focuses on global structural information, whereas later timesteps emphasize high-frequency details. Moreover, as verified in the Draft (Clark et al., 2023), optimizing at the final timestep also drives passive optimization of preceding timesteps.
@@ B.2 KDE of Cosine Similarity Distributions in Deep Feature Space
B.3 Training details
We select two representative Diffusion-SR paradigms and conduct experiments across models of different scales: DiT4SR and ResShift. Specifically, DiT4SR is built upon Stable Diffusion 3.5 with 2716.8M parameters, which shares a similar architecture with Stable Diffusion 3 (Esser et al., 2024). The training process is conducted on 512 × 512 resolution images. We train our model with a constant learning rate of $1e^{-6}$ with a batch size of 8. During inference, we adopt the default sampling schedule of SD3.5 with 40 sampling steps (T). The scale of classifier-free guidance (CFG) is set to 8 in our experiments. Following (Duan et al., 2025), the prompt of the input LR image is obtained from LLaVA (Liu et al., 2023).
ResShift is a lightweight model trained from scratch with 16.7M parameters, which shares a similar architecture with UNet (Ho et al., 2020), the training process is conducted on 256 × 256 resolution images. We train our model with a constant learning rate of $1e^{-6}$ with a batch size of 8. During inference, we adopt the default sampling schedule of ResShift with 15 sampling steps (T).
All the evaluation metrics are implemented by PyIQA (Chen and Mo, 2022). Note that the metric of ‘ClipIQA’ is implemented with the setting of ‘clipiqa+_vitL14 512’ provided by PyIQA.
@@ B.3 Training details
Appendix C Extended Ablation
C.1 Ablation on Loss Components
@@
As presented in Table 5 and Table 6, utilizing solely the reward signal without additional constraints leads to a significant improvement in perceptual metrics. However, this gain is achieved at the expense of severe structural distortion, indicating that the model has succumbed to reward hacking. The introduction of either semantic or structural constraints results in a regression in perceptual scores but successfully mitigates distortion. Ultimately, by integrating both constraints, our model achieves a stable and synergistic improvement across both perceptual quality and distortion metrics.
C.2 Ablation on Effectiveness of Bidirectional Reward-Guided Diffusion
@@ C.3 Dynamic Weight
Model
$\mathcal{L}_{\text{struct}}$
$\mathcal{L}_{\text{sem-align}}$
$\mathcal{L}_{\text{reward}}$
ClipIQA $\uparrow$ LPIPS $\downarrow$
A 0.541 0.355
B 0.538 0.338
C 0.544 0.326
FULL 0.565 0.322
@@
Appendix D More visual results
@@
Appendix E User Study Configuration
@@
Appendix F Related Methods
Another line of research involves $D^{2}PO$ (Wu et al., 2025), which explores preference optimization via synthetic data rollouts. Although $D^{2}PO$ provides a robust baseline for aesthetic alignment, its application to Real-World ISR is limited by the neglect of LR-to-HR distribution gaps. Our Bird-SR framework treats this distribution shift as a primary constraint. Notably, our approach is orthogonal to $D^{2}PO$; while our Bird-SR provides the base preference learning objective, $D^{2}PO$’s HPO (Hierarchical Preference Optimization) and hybrid rewards can be seamlessly integrated into our Bird-SR framework to handle complex real-world degradations.
Appendix G Limitations & Future Work.
While Bird-SR demonstrates strong performance on real-world super-resolution, its effectiveness is inherently tied to the quality of the reward model. The current reward formulation primarily captures global perceptual preferences and may be insufficient for modeling fine-grained, region-specific visual attributes such as local textures, edges, or small structures. Future work could explore more precise and fine-grained reward models, for example by incorporating spatially adaptive or multi-scale reward signals, as well as learning region-aware preferences from richer human feedback. Such extensions may further enhance the robustness and controllability of reward-guided diffusion for real-world image restoration.
Appendix H Broader Impacts

### v2

@@ 5.1 Experimental Settings
Datasets. We adopt a combination of images from DIV2K [div2k], DIV8K [div8k], Flickr2K [flickr2k], and the first 10K face images from FFHQ [ffhq] during training. Flickr8K [Flickr8K] is used for real-world low-resolution dataset. The degradation pipeline of RealESRGAN [realesrgan] is utilized to synthesize LR-HR training pairs with the same parameter configuration as baseline methods. Following DiT4SR [DiT4SR], the resolutions are set to 128 × 128 and 512 × 512 for LR and HR images, respectively. Following ResShift [ResShift], the resolutions are set to 64 × 64 and 256 × 256 for LR and HR images.
Since our method specifically focuses on Real-ISR task, we evaluate our model on four widely used real-world datasets, including DrealSR [drealsr], RealSR [realsr], RealLR200 [SeeSR], and RealLQ250 [DreamClear]. All experiments are conducted with the scaling factor of ×4. DrealSR and RealSR respectively consist of 93 and 100 images. Following DiT4SR, center-cropping is adopted for these two datasets, and the resolution of LR images is set to 128 × 128.
Metrics. Following [DiT4SR, DreamClear, SUPIR], PSNR and SSIM [ssim] are inadequate for measuring perceptual differences. Thus, most studies use the perceptual measurement LPIPS [lpips] for image fidelity. We use MUSIQ [musiq], MANIQA [maniqa], ClipIQA [ClipIQA], and LIQE [liqe] as non-reference metrics to measure image quality. In addition, we conduct a user study to comprehensively assess both fidelity and perceptual quality.
@@
5.2 Comparison with Other Methods
We compare our method with state-of-the-art RealISR methods, including diffusion-based approaches (i.e., ResShift [ResShift], StableSR [StableSR], SeeSR [SeeSR], DiffBIR [DiffBIR], and SUPIR [SUPIR]), as well as diffusion-based methods with DiT architecture (i.e. DreamClear [DreamClear] and DiT4SR [DiT4SR]). As shown in Table 1, our method achieves strong quantitative performance across all four real-world super-resolution benchmarks. Additional visual results, including the outputs of ResShift, are presented in the Appendix. As illustrated in Figure 4, our approach produces more realistic and natural reconstructions in real-world scenarios, with clearer and more faithful details while effectively suppressing artifacts. These results indicate that our method effectively mitigates the impact of input distribution shift and improves perceptual quality on real-world low-resolution images. By leveraging bidirectional reward feedback learning and jointly training on synthetic and real low-resolution data, our method demonstrates superior performance and robustness, all without adding any computational cost or latency during inference.
5.3 Ablation Experiments
To further demonstrate the effectiveness of each component, we conduct the ablation study on RealSR [realsr] with MUSIQ and LPIPS as evaluation metrics. All variants are trained using the same settings as the full model for fair comparison. Additional details are provided in the Appendix.
@@ 5.3 Ablation Experiments
Table 2: User study results on real-world datasets. The percentages denote the frequency with which Ours was preferred over each compared approach (realism and fidelity).
@@ 5.3.1 Effectiveness of Bidirectional Reward-Guided Diffusion
4. Mixed forward and reverse: The full bidirectional optimization framework that jointly leverages synthetic paired data and real-world low-resolution data.
@@
Setting MUSIQ $\uparrow$ LPIPS $\downarrow$
@@
5.3.2 Dynamic Weights.
We investigate the impact of different distortion-perception weighting strategies across diffusion timesteps on model performance. Specifically, we compare several weighting schedules for combining distortion loss $\mathcal{L}_{\text{struct}}$ and reward-based loss $\mathcal{L}_{\text{reward}}$ in our bidirectional reward-guided SR framework: $\lambda(t)=\left(\frac{t}{T}\right)^{\gamma}$, which gradually shifts from distortion to perceptual reward, where $\gamma>0$ controls the decay rate.
Analysis. Table 4 summarizes the quantitative results, and Figure 6 visualizes the different weighting schedules. Our ablation study on power-law weighting with varying exponents reveals a critical trade-off [8578750, RFSR, 10.1007/978-3-030-58621-8_37]: stronger structural constraints at early timesteps reduce distortion but limits perceptual reward gains. Conversely, insufficient early-stage constraints lead to structural degradation, which triggers reward hacking. By identifying an optimal weighting scheme, we enhance perceptual preference while preserving structural integrity.
@@
Table 4: Ablation study on different distortion–perception weighting. We report perceptual quality of MUSIQ and LPIPS.
5.3.3 Ablation on Loss Components.
To verify the effectiveness of our proposed structural constraint and semantic alignment, we conduct ablation studies on DiT4SR. For structural constraint, we employ perceptual supervision via LPIPS loss between the policy model’s output and the ground truth. We deliberately eschew conventional $L_{1}$ or $L_{2}$ constraints, as they tend to induce over-smoothed results. For semantic alignment, we leverage DINO in the feature space to align the outputs of the reference and policy models. Since DINO captures high-level semantics, it effectively constrains structural distortions arising from reward hacking while reserving optimization overhead for pixel-level detail preferences. This dual-pronged approach ensures both the stability and efficacy of the training process.
Analysis. As presented in Table 5, utilizing solely the reward signal without additional constraints leads to a significant improvement in perceptual metrics. However, this gain is achieved at the expense of severe structural distortion, indicating that the model has succumbed to reward hacking. The introduction of either semantic or structural constraints results in a regression in perceptual scores but successfully mitigates distortion. Ultimately, by integrating both constraints, our model achieves a stable and synergistic improvement across both perceptual quality and distortion metrics.
Model
$\mathcal{L}_{\text{struct}}$
$\mathcal{L}_{\text{sem-align}}$
$\mathcal{L}_{\text{reward}}$
ClipIQA $\uparrow$ LPIPS $\downarrow$
A 0.541 0.355
B 0.538 0.338
C 0.544 0.326
FULL 0.565 0.322
6 Limitations & Future Work.
While Bird-SR demonstrates strong performance on real-world super-resolution, its effectiveness is inherently tied to the quality of the reward model. The current reward formulation primarily captures global perceptual preferences and may be insufficient for modeling fine-grained, region-specific visual attributes such as local textures, edges, or small structures. Future work could explore more precise and fine-grained reward models, for example by incorporating spatially adaptive or multi-scale reward signals, as well as learning region-aware preferences from richer human feedback. Such extensions may further enhance the robustness of reward-guided diffusion for real-world image restoration.
@@ 7 Conclusion
8 Detailed Formulations of Diffusion-based Super-Resolution
@@
8.1 Diffusion Models Trained from Scratch for Super-Resolution
The first category directly trains a diffusion model for super resolution using paired low-resolution (LR) and high-resolution (HR) data. Representative methods such as ResShift formulate the forward diffusion process as an interpolation between the HR image $x_{0}$, the LR-conditioned signal $x_{1}$, and Gaussian noise.
@@ 8.1 Diffusion Models Trained from Scratch for Super-Resolution
8.2 Pretrained Diffusion Models with Conditional Adaptation
@@
A common approach is to introduce an external conditioning module, such as ControlNet, which injects LR structural information into intermediate layers of the frozen pretrained diffusion backbone. Alternatively, lightweight parameter-efficient tuning strategies, such as Low-Rank Adaptation [hu2022lora] (LoRA), are employed to fine-tune a subset of model parameters while preserving the pretrained generative prior. Specifically, the forward diffusion process remains unchanged and is given by:
@@
where $x_{0}$ denotes the target high-resolution image. The low resolution observation $y$ is not involved in the forward noising process but is provided as a conditioning signal to the denoising network. Formally, the reverse diffusion process can be written as:
@@
9 Implementation details
9.1 Local Binary Pattern (LBP) Texture Features
@@
9.2 KDE of Cosine Similarity Distributions in Deep Feature Space
We further analyze the relationship between low-resolution and high-resolution (HR) image pairs in deep feature space by computing the cosine similarity between their feature embeddings extracted from a pretrained network. Kernel density estimation (KDE) is employed to estimate the distribution of cosine similarities across the dataset. Across different frequency bands, synthetic data exhibits a notable distribution expansion coupled with a significant domain shift, as shown in Figure 9. In the low-frequency band, synthetic LR–GT pairs maintain high cosine similarity concentrated near 1.0, closely matching real-world distributions and preserving global structures. However, as the frequency increases, the synthetic distribution becomes increasingly broader and shifts toward lower similarity values; specifically, the mid-frequency band reveals discrepancies in meso-scale textures, while the high-frequency band displays the most pronounced flattening and leftward shift. This trend suggests that while synthetic degradations cover a wider range of variations, they fail to accurately replicate the complex, fine-grained artifacts of real-world processes, identifying high-frequency components as the primary driver of the overall input distribution shift. Therefore, when performing reward feedback learning on real-world data, we primarily apply optimization at the final timestep. This is because, during early timesteps, the diffusion model focuses on global structural information, whereas later timesteps emphasize high-frequency details. Moreover, as verified in the Draft [draft], optimizing at the final timestep also drives passive optimization of preceding timesteps.
@@ 9.2 KDE of Cosine Similarity Distributions in Deep Feature Space
9.3 Training details
We select two representative Diffusion-SR paradigms and conduct experiments across models of different scales: DiT4SR and ResShift. Specifically, DiT4SR is built upon Stable Diffusion 3.5 with 2716.8M parameters, which shares a similar architecture with Stable Diffusion 3 [sd3]. The training process is conducted on 512 × 512 resolution images. We train our model with a constant learning rate of $1e^{-6}$ with a batch size of 8. During inference, we adopt the default sampling schedule of SD3.5 with 40 sampling steps (T). The scale of classifier-free guidance (CFG) is set to 8 in our experiments. Following [DiT4SR], the prompt of the input LR image is obtained from LLaVA [liu2023llava]. For DiT4SR, we update only the control network while keeping all other components frozen. For ResShift, we freeze the first $70\%$ of the network parameters to retain its pre-trained super-resolution capability, and only update the remaining $30\%$ during training.
ResShift is a lightweight model trained from scratch with 16.7M parameters, which shares a similar architecture with UNet [DDPM], the training process is conducted on 256 × 256 resolution images. We train our model with a constant learning rate of $1e^{-6}$ with a batch size of 8. During inference, we adopt the default sampling schedule of ResShift with 15 sampling steps (T).
All the evaluation metrics are implemented by PyIQA [pyiqa]. Note that the metric of ‘ClipIQA’ is implemented with the setting of ‘clipiqa+_vitL14 512’ provided by PyIQA.
@@ 9.3 Training details
10 Extended Ablation
10.1 Ablation on Loss Components
@@
As presented in Table 6, utilizing solely the reward signal without additional constraints leads to a significant improvement in perceptual metrics. However, this gain is achieved at the expense of severe structural distortion, indicating that the model has succumbed to reward hacking. The introduction of either semantic or structural constraints results in a regression in perceptual scores but successfully mitigates distortion. Ultimately, by integrating both constraints, our model achieves a stable and synergistic improvement across both perceptual quality and distortion metrics.
10.2 Ablation on Effectiveness of Bidirectional Reward-Guided Diffusion
@@ 10.3 Dynamic Weight
11 More visual results
@@
12 User Study Configuration
@@
13 Related Methods
Another line of research involves $D^{2}PO$ [wu2025dposr], which explores preference optimization via synthetic data rollouts. Although $D^{2}PO$ provides a robust baseline for aesthetic alignment, its application to Real-World ISR is limited by the neglect of LR-to-HR distribution gaps. Our Bird-SR framework treats this distribution shift as a primary constraint. Notably, our approach is orthogonal to $D^{2}PO$; while our Bird-SR provides the base preference learning objective, $D^{2}PO$’s HPO (Hierarchical Preference Optimization) and hybrid rewards can be seamlessly integrated into our Bird-SR framework to handle complex real-world degradations.
14 Broader Impacts